OpenStudy (anonymous):

9.02] What are the solutions of x2 + 8x − 20 = 0? x = 2, x = 10 x = −2, x = −10 x = 2, x = −10 x = −2, x = 10

OpenStudy (anonymous):

The answer would be the second to last one because -2 * 10 = -20 ans -2+10 = 8 put that into brackets as (x-2) (x+10)..... so x=2 as 2-2 =0, x=-10 as -10+10 = 0

OpenStudy (anonymous):

Oh, okay, thx! thx for the explanation!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Could you help me with one more?

OpenStudy (anonymous):

[9.01] What is the vertex for the graph of y = −2x2 + 8x − 13? (2, 11) (2, −5) (−2, −37) (−2, −21)

OpenStudy (anonymous):

I'm not too sure about this one it requires a graph and i cant picture it

OpenStudy (anonymous):

Oh, okay, then...this one?

OpenStudy (anonymous):

9.01] When the quadratic equation y = (x − 4)(2x + 5) is written in standard form, what is the value of the coefficient "b"? 2 3 −3 −20 is it a?

OpenStudy (anonymous):

this one you have to expand the brackets so it becomes, 2x^2 - 8x - 20 + 5x simplify it to be ... 2x^2 - 3x -20 so the answer would be -3

OpenStudy (anonymous):

For the second question: y = −2x2 + 8x − 13 = -(x-2)^2- 5 Therefore, vertex will be (2, -5).

OpenStudy (anonymous):

Sorry typo .... For the second question: y = −2x2 + 8x − 13 = -2(x-2)^2- 5 Therefore, vertex will be (2, -5).

OpenStudy (anonymous):

could you show how you got that without the graph, i would like to know

OpenStudy (anonymous):

We have an equation ax2 + bx + c =0 The vertex is \[V=(\frac{ -b }{ 2a },c-\frac{ b^{2} }{ 4a })\]

OpenStudy (anonymous):

Hint: perfect squere y = −2x2 + 8x − 13 = -2(x-2)^2- 5 = 0 x-2 = 0, x =2 y = -5 V = (2, -5)

OpenStudy (anonymous):

OMG, thank you so much! LOL, it was so nice of you to help!